Write True (T) or False ( F ) against each of the following statements :
 16 : 24 :: 20 : 30

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

16: 24 :: 20: 30
16 / 24 = 2 / 3
20 / 30 = 2 / 3
Hence, 16: 24 = 20: 30
Therefore True

Thanks